OpenStudy (anonymous):

16. Find the balance in the account. $2,550 at 3% compound interest for 18 months (1 point) HINT: Use the compounded Interest Formula: B = P( 1 + r )t B is the final amount including the principal. P is the principal amount (initial or starting amount). r is the rate of interest per year (convert it to a decimal). t is the number of years invested. • $2,665.61 • $2,665.61 • $2,689.35

OpenStudy (anonymous):

Fv=p(1+r)^n FV=2550(1+0.03)^1.5=$2665.61

OpenStudy (anonymous):

Thank you so much!